# BConformeR: A Conformer Based on Mutual Sampling for Unified Prediction of Continuous and Discontinuous Antibody Binding Sites

### v2

@@ Abstract
Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and for advancing our understanding of immune responses. Despite in silico methods that have been proposed to predict both linear (continuous) and conformational (discontinuous) epitopes, they consistently underperform in predicting conformational epitopes. In this work, we propose a conformer-based model trained on antigen sequences derived from 1,080 antigen-antibody complexes, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences. Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes. Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on both linear and conformational epitopes. The data and code are available at https://github.com/Jerry-alt19/BConformeR.
@@ 1 Introduction
Figure 1: Overview of epitope classification and BConformeR prediction. (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388). Data were derived from 1,432 antigen records (Appendix C.2). (B) Epitope prediction workflow. Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
@@
Figure 2: BConformeR overview. (A) Preparation pipeline of the data. (Right) 1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes. (Left) Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280. (B) Conformer architecture. Stem module applies convolution and max-pooling to the original ESM-2 embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; Feature coupling units (FCU-Up and FCU-Down) connect the Conv and Trans blocks, with 11 FCU layers in total; The classifier linearly fuses features from both Conv and Trans branches.
@@
Here, we propose BConformeR, a Conformer-based model for continuous and discontinuous BCR epitope prediction. Antigen sequences from 1,180 Ag-Ab complexes were embedded using ESM-2, with 1,080 used for training and 100 for evaluation. In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on both linear and conformational epitopes. To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
2 Related Work
2.1 Antibody-agnostic prediction models
Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction (Appendix A). SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0 . BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0 . Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D ; AdaBoost . ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet . SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed using an ensemble of five pre-trained Structure-aware Protein language models (SaProt) SEMA2.0 ; SaProt . DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0 ; ESM-IF1 .
However, existing models rarely incorporate both local features and long-range dependencies within antigen sequences, showing limited performance on conformational epitopes. Therefore, we adopted a Conformer based on mutual sampling, employing CNNs to extract local features and Transformers to capture long-range dependencies within antigen sequences.
2.2 Antibody-aware evaluation models
Despite advances in epitope prediction, evaluating Ag-Ab binding confidence remains challenging, which is crucial for vaccine selection. AbEpitope-1.0, a recent method for predicting antibody targets by use of AlphaFold and inverse folding, significantly improving the prediction accuracy of Ag-Ab interfaces. They constructed negative Ag-Ab samples (false binding pairs) by swapping antibody chains across real complexes in the dataset. Each sample was labeled as either true or false to train the model to distinguish authentic Ag-Ab interactions from decoys. The final model, AbEpiTarget-1.0, achieved superior performance in antibody-sensitive scoring for target prediction AbEpitope .
In this work, we utilize the real Ag-Ab complex dataset provided by AbEpitope-1.0, employing its epitope identification pipeline to ensure the data quality (Fig. 2A).
@@ 3 Methods
3.1 Epitope Extraction
3.1.1 Epitope identification
During immune recognition, epitopes are typically composed of multiple spatially adjacent residues. To capture this interaction, the notions of residue-neighbor and residue-patch were introduced. A residue-neighbor is defined when the minimum distance between any atoms of two residues is less than 4 Å, and a residue-patch refers to a group of residues whose atoms lie within 10 Å of a central residue epitope_id . In our study, we focused on the residue-neighbor relationship to capture the tight spatial interactions, which is given as:
$\min_{a\in r_{i},\,b\in r_{j}}\|a-b\|<4\,\text{\AA }$, (1)
where $r_{i}$ and $r_{j}$ are residues from the antigen and antibody chains respectively, and $a$, $b$ represent the heavy atoms within these residues.
3.1.2 Epitope classification
Generally, epitopes are separated into multiple Regions (R) and Fragments (F) for each antigen sequence. Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues. Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl (Appendix B). Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig. 1A).
@@ 3.1.2 Epitope classification
stage output CNN Branch FCU Transformer Branch
stem 1024 7, 320
3 max pooling
s1 1024 $\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$ - 5, 1536 $\times$1
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$ $\times$3
$1,1536]\longrightarrow$
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$
$\longleftarrow[1,160]$
s2 1024 $\left[\begin{array}[]{l}\text{1, 320}\\[-1.00006pt] \text{3, 320}\\[-1.00006pt] \text{1, 1280}\end{array}\right]$ $\times$4
$1,1536]\longrightarrow$
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 320}\\[-1.00006pt] \text{3, 320}\\[-1.00006pt] \text{1, 1280}\end{array}\right]$
$\longleftarrow[1,320]$
s3 1024 $\left[\begin{array}[]{l}\text{1, 640}\\[-1.00006pt] \text{3, 640}\\[-1.00006pt] \text{1, 2560}\end{array}\right]$ $\times$4
$1,1536]\longrightarrow$
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 640}\\[-1.00006pt] \text{3, 640}\\[-1.00006pt] \text{1, 2560}\end{array}\right]$
$\longleftarrow[1,640]$
[…]
$\alpha\times\mathrm{Conv}_{\mathrm{cls}}+(1-\alpha)\times\mathrm{Trans}_{\mathrm{cls}}$
Table 1: Architecture of BConformeR (BConformer-12), a 1D adaptation of Conformer Conformer21 . The MHSA-12 indicates multi-head self-attention with 12 heads operating on 1536-dimensional embeddings (128 dimensions per head). $\mathrm{Conv}_{\mathrm{cls}}$ denotes the 1D convolutional classifier operating on CNN features, and $\mathrm{Trans}_{\mathrm{cls}}$ denotes the linear classifier operating on Transformer outputs. The final prediction is obtained by a weighted fusion of the two.
@@ 3.2.1 Overview
To take advantage of both local and global representations for epitope prediction, we construct BConformeR, a dual-branch architecture adapted from Conformer Conformer21 , as shown in Table 1. BConformeR consists of a shared convolutional stem, a CNN branch, a Transformer branch, Feature Coupling Units (FCUs) for bidirectional feature exchange, and two per-position classifiers (Fig. 2B). Since our task is per-residue epitope prediction, all antigen sequences are padded to a fixed length of $L$ (e.g., 1024) and this length is preserved throughout the model.
The stem consists of a 1D convolution (kernel size 7) and a 1D max pooling (kernel size 3), designed to capture contextual features for each residue. And the embedding dimension $C$ (e.g., 1280) is reduced to $C/4$ channels after the stem. In the first block, both branches process identical initial features and therefore no feature coupling is applied. From the second block onward, FCUs are inserted to progressively fuse local and global representations. The input is processed by alternating CNN and Transformer blocks, where at each stage, CNN-derived local features are passed to the Transformer branch, and Transformer-derived global representations are fed back to the CNN branch. The CNN branch gradually expands to $2C$ channels, and the Transformer branch maintains $D$-dimensional (e.g., 1536) token embeddings at the end.
@@ 3.2.2 CNN branch
The CNN branch is composed of three stages after the stem, each consisting of repeated convolutional blocks based on the ResNet bottleneck design ResNet . Given an input feature sequence $\mathbf{x}\in\mathbb{R}^{C^{\prime}\times L}$, a bottleneck performs three sequential convolutions with a residual connection applied to $\mathbf{x}$:
@@
The channel dimension is first downsampled, then expanded by a factor of 4 within each bottleneck. The first convolution block contains a single bottleneck, while subsequent blocks include two bottlenecks: one applied before the Transformer branch, and the other afterward. Since convolution kernels slide over representation sequences with overlap, the CNN blocks are well-suited to capture residue-level contextual information, providing local features for the Transformer branch.
@@ 3.2.3 Transformer branch
The Transformer branch processes residue embeddings using 12 repeated Transformer blocks. Each block follows the standard architecture introduced in ViT ViT , but adapted to 1D protein sequences. As shown in Table 1, each block contains a multi-head self-attention module and a two-layer feed-forward MLP block. Both submodules are preceded by Layer Normalization (LN) with residual connections. Specifically, given the input sequence $\mathbf{x}\in\mathbb{R}^{L\times D}$, a single Transformer block updates the representation as follows:
@@
where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads Attention , and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$. The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
@@ 3.2.4 Feature coupling units
To enable bidirectional interaction between the local CNN features and global Transformer representations, we introduce FCUs starting from the second block. FCUs consist of two modules: FCU-Down, which passes CNN features to the Transformer; and FCU-Up, which feeds Transformer outputs back to the CNN.
FCU-Down projects the CNN features $\mathbf{F}_{\mathrm{CNN}}\in\mathbb{R}^{C_{1}\times L}$ into the Transformer embedding space $\mathbb{R}^{L\times D}$ via a $1\times 1$ convolution and batch normalization, followed by transposition to align the dimensions. The projected features $\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}$ are then added to the Transformer input $\mathbf{F}_{\mathrm{Trans}}$:
@@
FCU-Up maps the Transformer features $\widetilde{\mathbf{F}}_{\mathrm{Trans}}\in\mathbb{R}^{L\times D}$ back to the CNN feature space $\mathbb{R}^{C_{2}\times L}$ by transposing first, then applying a $1\times 1$ convolution and batch normalization. The transformed features $\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}$ are added to the original CNN features $\mathbf{F}_{\mathrm{CNN}}$:
@@
3.3 Performance Metric
3.3.1 AgIoU
To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such (Appendix D):
$\text{AgIoU}=\frac{\text{Res}_{\text{true}}\cap\text{Res}_{\text{pred}}}{\text{Res}_{\text{true}}\cup\text{Res}_{\text{pred}}}.$ (8)
@@ 4.1.1 Dataset
We collected 1,300 Ag–Ab complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0. Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures (Appendix C.1). For each complex, we selected the top-ranked model and excluded those with a pLDDT score below 0.82. Interface residues were extracted based on a distance cutoff of 4 Å between heavy atoms on the antigen and antibody chains (1). To enhance data quality and limit parameter overhead, we removed complexes containing insufficient epitope residues or antigens exceeding 1024 residues in length. Finally, the dataset contains 1,180 Ag-Ab complexes in total: 1,080 used for training BConformeR, and 100 set aside for evaluation and ablation. And the antigen chains annotated with epitope labels serve as the input.
Figure 3: Improved performance on epitope prediction. ROC-AUC and PR-AUC curves plot on the blind test set of 24 antigen records. (A) ROC-AUC curves. (B) PR-AUC curves.
Method Threshold F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$ PCC $\uparrow$ AgIoU $\uparrow$
SEPPA-3.0 0.100 ± 0.020 0.158 ± 0.009 0.237 ± 0.025 0.058 ± 0.014 0.067 0.088 ± 0.005
BepiPred-3.0 0.220 ± 0.010 0.276 ± 0.007 0.209 ± 0.011 0.194 ± 0.003 0.215 0.155 ± 0.003
SEMA-1D 2.0 0.500 ± 0.030 0.270 ± 0.005 0.213 ± 0.006 0.188 ± 0.002 0.217 0.147 ± 0.002
DiscoTope-3.0 0.300 ± 0.020 0.303 ± 0.007 0.231 ± 0.004 0.227 ± 0.007 0.276 0.167 ± 0.002
BConformeR 0.330 ± 0.020 0.349 ± 0.004 0.262 ± 0.003 0.296 ± 0.003 0.347 0.203 ± 0.002

### v3

@@ Abstract
Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and advancing our understanding of immune responses. Despite in silico methods that have been proposed to predict both linear (continuous) and conformational (discontinuous) epitopes, they consistently underperform in predicting conformational epitopes. In this work, we propose Conformer-based models trained separately on AlphaFold-predicted structures and experimentally determined structures, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences. Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes. Experimental results show that our model outperforms existing baselines in terms of MCC, ROC-AUC, PR-AUC, and F1 scores on both linear and conformational epitopes. Code is available at https://anonymous.4open.science/r/bconformer-epitope-519D.
@@ 1 Introduction
Figure 1: Overview of epitope classification and BConformeR prediction. (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388). Data were derived from 1,432 antigen records. (B) Epitope prediction workflow. Antigen inputs are processed to predict linear and conformational epitopes, which enables applications in vaccine design and diagnostic development.
@@
Figure 2: BConformeR overview. (A) Preparation pipeline. (Right) Preparation of Alphafold-predicted structures and epitope extraction. (Left) Antigen embeddings. (B) Architecture overview. Stem module applies convolution and max-pooling to the ESM Cambrian (ESMC) embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; FCU-Up and FCU-Down connect the Conv and Trans blocks; The classifier linearly fuses features from both branches.
@@
Here, we propose BConformeR, a Conformer-based model for continuous and discontinuous BCR epitope prediction. BConformeR outperforms baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including MCC, ROC-AUC, PR-AUC, and F1 score on both linear and conformational epitopes. To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
@@ 2 Methods
2.1 Epitope classification
To extract epitope residues, we define the residue–neighbor, where an antigen residue and an antibody residue are considered binding sites if the minimum distance between any pair of their heavy atoms is less than 4 Å epitope_id. Generally, epitopes are separated into multiple Regions (R) and Fragments (F) for each antigen sequence. Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues. Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl. Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig. 1A).
@@ 2.1 Epitope classification
  stage 1 stage 2 stage 3   CNN $\left[\begin{array}[]{l}1,144\\[-1.99997pt] 3,144\\[-1.99997pt] 1,576\end{array}\right]$ $\left[\begin{array}[]{l}1,144\\[-1.99997pt] 3,144\\[-1.99997pt] 1,576\end{array}\right]$ $\left[\begin{array}[]{l}1,144\\[-1.99997pt] 3,144\\[-1.99997pt] 1,576\end{array}\right]$ $\left[\begin{array}[]{l}1,288\\[-1.99997pt] 3,288\\[-1.99997pt] 1,1152\end{array}\right]$ $\left[\begin{array}[]{l}1,288\\[-1.99997pt] 3,288\\[-1.99997pt] 1,1152\end{array}\right]$ $\left[\begin{array}[]{l}1,576\\[-1.99997pt] 3,576\\[-1.99997pt] 1,2304\end{array}\right]$ $\left[\begin{array}[]{l}1,576\\[-1.99997pt] 3,576\\[-1.99997pt] 1,2304\end{array}\right]$   FCU - $\begin{array}[]{c}[1,1536]\\[-1.99997pt]$ $\big\downarrow\end{array}$ $\begin{array}[]{c}\big\uparrow\\[-1.99997pt] [1,144]\end{array}$ $\begin{array}[]{c}[1,1536]\\[-1.99997pt]$ $\big\downarrow\end{array}$ $\begin{array}[]{c}\big\uparrow\\[-1.99997pt] [1,288]\end{array}$ $\begin{array}[]{c}[1,1536]\\[-1.99997pt]$ $\big\downarrow\end{array}$ $\begin{array}[]{c}\big\uparrow\\[-1.99997pt] [1,576]\end{array}$   Trans $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$ $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$ $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$ $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$   $\times 1$ $\times 3$ $\times 4$ $\times 4$
Table 1: Interaction layers of BConformeR. MHSA denotes multi-head self-attention.
@@ 2.2.1 Overview
To take advantage of both local and global representations for epitope prediction, we construct BConformeR, a dual-branch architecture adapted from Conformer Conformer21. BConformeR consists of a shared convolutional stem, a CNN branch, a Transformer branch, Feature Coupling Units (FCUs) for bidirectional feature exchange, and two per-position classifiers (Fig. 2B).
Sequences are padded to match the longest length in each batch. The stem consists of a 1D convolution and a 1D max pooling, with the embedding dimension $C$ (e.g., 1152 generated by ESMC-600M) reduced to $C/4$ channels after the stem. Then as shown in Table 1, both branches process initial features and no feature coupling is applied in the first block. From the second block onward, FCUs are inserted to progressively fuse local and global representations. At each stage, CNN-derived features are passed to the Transformer branch, and Transformer-derived global representations are fed back to the CNN branch. The CNN branch gradually expands to $2C$ channels, and the Transformer branch maintains $D$-dimensional (e.g., 1536) token embeddings at the end.
@@ 2.2.2 CNN branch
The CNN branch is composed of three stages after the stem, each consisting of repeated convolutional blocks based on the ResNet bottleneck design ResNet. As shown in Table 1, given an input $\mathbf{x}\in\mathbb{R}^{C^{\prime}\times L}$, the bottleneck performs three sequential convolutions with a residual connection applied to $\mathbf{x}$:
@@
The channel dimension is first downsampled, then expanded by a factor of 4 within each bottleneck. The first convolution block contains a single bottleneck, while subsequent blocks include bottlenecks before and after the Transformer branch.
@@ 2.2.3 Transformer branch
The Transformer branch processes embeddings using 12 repeated Transformer blocks, with each block following a structure similar to ViT ViT. As shown in Table 1, given the input $\mathbf{x}\in\mathbb{R}^{L\times D}$, a single Transformer block updates the representation as follows:
@@
where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 24) heads Attention, and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$. The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
@@ 2.2.4 Feature coupling units
To enable bidirectional interaction between local CNN features and global Transformer representations, we introduce feature coupling units (FCUs) from the second block. FCU-Down projects CNN features into the Transformer space and adds them to the Transformer input:
@@
FCU-Up maps Transformer features back to the CNN space and merges them:
@@
2.3 Calibrated Score
Scores of residues in each antigen are normalized using the following formula:
$z_{i}=\frac{s_{i}-\mu_{\text{antigen}}}{\sigma_{\text{antigen}}},$ (7)
where $s_{i}$ denotes the raw score of residue $i$, and $\mu_{\text{antigen}}$ and $\sigma_{\text{antigen}}$ represent the mean and standard deviation of scores for that antigen, respectively. Both $\mu_{\text{antigen}}$ and $\sigma_{\text{antigen}}$ are estimated using two separate generalized additive models (GAMs) GAM, with one predicting the mean residue score from antigen length and the other predicting the standard deviation from the mean residue score DiscoTope3.0.
@@ 3.1.1 Dataset
We trained two models separately using experimentally solved and AlphaFold-predicted structures. The dataset of experimentally solved Ag–Ab complexes was obtained from AACDB AACDB_data. Redundant complexes were removed by retaining only one complex for each PDB ID, resulting in a curated dataset of 3,674 Ag-Ab complexes for training. For the AlphaFold-based dataset, we collected 1,300 Ag–Ab complex FASTA files from AbEpiTope-1.0 AbEpitope and modeled each complex using AlphaFold-2.3 ColabFold AFMultimer. Thirty structural predictions were generated for each complex, producing a total of 39,000 structures. For each complex, we retained the top-ranked model and discarded those with a pLDDT score below 0.82. After extracting epitopes for each complex, we further excluded those with insufficient epitope residues. The final dataset contained 1,180 Ag–Ab complexes, of which 1,080 were used for training BConformeR and 100 for validation and ablation studies.
Figure 3: Improved performance of BConformeR. (A–B) GAMs for estimating $\mu$ and std. (C–D) Comparison between raw and calibrated scores on the SARS-CoV-2 test set. (E–F) ROC-AUC and PR-AUC; BConformeR trained on AlphaFold-predicted structures and evaluated on the blind test set of 24 antigens. (G) PR-AUC; BConformeR trained on experimentally solved structures and evaluated on the SARS-CoV-2 test set.
As the datasets used among methods are not consistent, we assembled a SARS-CoV-2 test set from CoV-AbDab CoV-AbDab for models trained on experimentally solved structures, and selected 24 antigen records from the SEMA-1D 2.0 test sets to construct a blind test set for AlphaFold-based models.
3.1.2 Calibrated score
After training, the $\mu_{\text{antigen}}$ and $\sigma_{\text{antigen}}$ were estimated from the predicted score distributions (Fig. 3A,B). The scores were then calibrated based on different antigen lengths, which effectively reduced false positives while improving the true positive rate (Fig. 3C,D).
Method Alphafold structures Solved structures
F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$ F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$
SEPPA-3.0 0.158 ± 0.009 0.237 ± 0.025 0.058 ± 0.014 - - -
BepiPred-3.0 0.276 ± 0.007 0.209 ± 0.011 0.194 ± 0.003 0.299 ± 0.009 0.087 ± 0.001 0.153 ± 0.005
SEMA-1D 2.0 0.270 ± 0.005 0.213 ± 0.006 0.188 ± 0.002 0.327 ± 0.008 0.106 ± 0.007 0.193 ± 0.004
DiscoTope-3.0 0.303 ± 0.007 0.231 ± 0.004 0.227 ± 0.007 0.396 ± 0.023 0.095 ± 0.007 0.307 ± 0.006
BConformeR 0.349 ± 0.004 0.262 ± 0.003 0.296 ± 0.003 0.405 ± 0.004 0.135 ± 0.014 0.350 ± 0.014